ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60, 62 Mohsen, Mohammadzadeh, Tarbiat Modarres University, Iran Spectral Approximation to the Likelihood for an Intrinsic Gaussian Random Field Intrinsic Gaussian random fields generated by conditional autoregressive models are considered. A spectral approximations to the log-likelihood functions of an intrinsic random field was given by Künsch (1987). A rigorous treatment is given of the comparison between the exact and spectral maximum likelihood functions as the domain of observations increases. Reference: H. R. Künsch, Intrinsic autoregresions and related models on the two dimensional lattice. Biometrika 74(3) (1987), 517524.